A car travels from place A to the place B at 20Km/hour and return at 30Km/hour. The average speed of the car for the whole journey.

Grade:11

5 Answers

harihara kumar
10 years ago

20+30   =25

  2

TANAYRAJ SINGH CHOUHAN
10 years ago

Total average speed is simply = Total distance/Total time 
Lets say, 
D = distance travelled by the car in EACH direction 
t1 = time spent on onward trip 
t2 = time spent on return trip 
Thus, the total distance travelled by the car = D+D= 2D 
And, by the formula, Speed = Distance/Time 
S1 = D/t1 => t1 = D/S1 
S2 = D/t2 => t2 = D/S2 
Total average speed = Total Distance/Total time = 2D/(t1+t2) = 2D/(D/S1+D/S2) = 2S1*S2/(S1+S2) 

Remember this general formula for a total average speed problems: 
Total average speed = 2S1*S2/(S1+S2) 

So in the question S1 = 20km/hr 

                         &S2 = 30km/hr

So av. speed = 2 x 20 x 30/20+30

                   = 24km/hr

PLEASE APPROVE MY ANSWER IF YOU LIKE IT BY CLICKING YES BELOW THE ANSWER

AND BEST OF LUCK!!!!!!!!!!!!!!!!!!!!! 

Manoj Modalavalasa Kumar
10 years ago

Average speed:24Kmph

RAJESH KUMAR
5 years ago
Formula of average speed is equal to 2v1v2 / v1 + v2 = 2 ×20×30 /20 +30 = 1200/ 50 = 24 km / h,when an object goes to a point B from A and returns back from point B to A,then formula to calculate average speed is 2v1v2 / v1 + v2
Anjali kumari
4 years ago
Let the distance covered from A to B is S.Then the distance covered from B to A is also S . Therefore, average speed= total distance travelled / total time taken . Now S+S/S/20+ S/30 =2S/5S/60 =2S×60/5S =24.